OpenStudy (anonymous):

how many solutions does the systens have? x= -4y+4 2x+8y=8

zepdrix (zepdrix):

\[\large x=-4y+4 \qquad \rightarrow \qquad x+4y=4\]\[\large 2x+8y=8 \quad \text{dividing by 2 gives us,}\quad x+4y=4\]See how they are the same equation if we alter the second one a little bit? They share every point in common. Meaning there are infinite solutions.